OpenStudy (anonymous):

Solve the following system of equations and show all work. y = 2x2 y = –3x −1

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

\[y=2.x^{2}\]?

OpenStudy (anonymous):

no y=2x^2

OpenStudy (triciaal):

set the y's equal and solve for x then substitute this value in one of the expressions for y to get the value of y

OpenStudy (mathmate):

Use comparison: y=2x^2 y=-3x-1 Since they are both y, they are equal, so put 2x^2=-3x-1, or transposing, solve 2x^2+3x+1=0 (by factoring) to solve the problem.

OpenStudy (triciaal):